OpenStudy (anonymous):

1.The water park says it has 2 times as many giant inner tubes as water sprinklers and 4 times as many sprinklers as water slides. The park has 16 water slides. a) How many giant inner tubes does the water park have? (b) Explain how you found your answer.

OpenStudy (anonymous):

so if water slides are w sprinklers would be s tubes would be t s=w*4 t=s*2=w*8 w+s+t=total w=12 s=48 t=96 So 96 tubes I showed work for (b)